# DAPS++: Rethinking Diffusion Inverse Problems with Decoupled Posterior Annealing

### v1

@@ 3.4 Connections to Existing Diffusion-Based Solvers
Remark. The E–M perspective introduced above also clarifies the behavior of existing diffusion-based solvers. Although many diffusion-based inverse problem approaches are presented as posterior samplers, recent studies show that their updates often resemble MAP-style iterative refinements [xu2025rethinking], which aligns with the interpretation adopted in this work. A representative example is diffusion posterior sampling (DPS) [chung2022diffusion], where $\mathbf{x}_{t}$ and $\mathbf{x}_{t+\Delta t}$ are linked through a combined update consisting of a diffusion prior term and a data-consistency term: $s_{\theta}^{*}(\mathbf{x}_{t},\sigma_{t})\;-\;\rho\,\nabla_{\mathbf{x}_{t}}\|\mathbf{y}-\mathcal{A}(\hat{\mathbf{x}}_{0}(\mathbf{x}_{t}))\|_{2}^{2}$, with $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})$ computed via Eq. 6.
From Eq. 3, the diffusion prior gradient and the likelihood gradient are statistically independent. This allows the DPS update to be rewritten as
$\mathbf{x}_{t-1}^{\prime}$ $=\mathbf{x}_{t}+\sigma_{t}^{2}\,s_{\theta}(\mathbf{x}_{t},\sigma_{t})+\sigma_{t-1}\mathbf{z}$ (10)
$\quad-\eta_{t}\Big[\nabla_{\mathbf{x}_{0}}\,\|\mathbf{y}-\mathcal{A}(\mathbf{x}_{0})\|_{2}^{2}\Big]_{\mathbf{x}_{0}=\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})}.$
where $\mathbf{z}\!\sim\!\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\eta_{t}$ denotes the step size.
Under our formulation, this update naturally decomposes into two decoupled operations:
  1. 1.
    Expectation step: estimate the clean mean $\hat{\mathbf{x}}_{0}$ from $\mathbf{x}_{t}$ using the diffusion score derived from Tweedie’s formulation;
  2. 2.
    Maximization step: refine this estimate through a single MAP-style likelihood update, followed by a re-noising operation that transitions the system to the next noise level $\sigma_{t-1}$.
This reinterpretation indicates that what is commonly described as posterior sampling in diffusion-based inverse problems is, in effect, a sequential E–M procedure alternating between diffusion-based estimation and data-driven optimization. By making this separation explicit, DAPS++ provides a clearer theoretical grounding and a more stable computational structure for general inverse problems, reducing reliance on explicit prior constraints and improving overall robustness.
@@ 4 Experiments
Figure 3: Qualitative results on representative inverse problems: (a) Gaussian blur reconstruction on the FFHQ-256 dataset compared with selected baselines. (b) Nonlinear blur reconstruction on ImageNet, comparing DAPS++-50 and DAPS++-100 alongside selected baselines. (c) High dynamic range (HDR) reconstruction on the FFHQ-256 dataset. (d) Motion blur reconstruction on the FFHQ-256 dataset. (e) Box inpainting results on ImageNet, where DAPS++ accurately restores structural details. (f) $8\times$ super-resolution using a pre-trained ImageNet diffusion model, demonstrating strong visual fidelity.
@@ 4.1 Experimental Setup
We evaluate our method using pixel-space pre-trained diffusion models on the FFHQ [chung2022diffusion] and ImageNet [dhariwal2021diffusion] datasets. For the noise schedule, we adopt the POLY(-7) discretization from EDM [karras2022elucidating]. To improve efficiency, $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})$ is generated using Eq. 6 while $\sigma_{t}>\bar{\sigma}$, and a single RK4 update is used once $\sigma_{t}\leq\bar{\sigma}$. Empirically, we set $\bar{\sigma}=10\times\gamma$ to balance computational cost and reconstruction accuracy. The total number of neural function evaluations (NFEs) remains low because each iteration requires only one network evaluation, further reduced by an annealed step-size schedule.
We use DAPS++-50 (50 annealing steps) for FFHQ and DAPS++-100 (100 steps) for ImageNet. Each DAPS++ iteration performs 4–8 MCMC refinement steps, except for nonlinear deblurring, which requires around 50 refinement steps in the M-step. Learning rates are tuned individually for each task. Additional details regarding model configurations, samplers, and hyperparameters appear in the Appendix, along with a discussion of sampling efficiency in Fig. 5.
@@ 4.2 Results
Method SR ×4 Inpainting Gaussian Blur Motion Blur Nonlinear Blur HDR
@@
DDNM 0.720 / 0.290 / 152.94 0.810 / 0.162 / 94.66 0.804 / 0.216 / 57.83
DDRM 0.820 / 0.191 / 76.07 0.792 / 0.210 / 88.90 0.786 / 0.218 / 89.40
DiffPIR 0.545 / 0.269 / 60.03 0.525 / 0.344 / 63.83 0.514 / 0.293 / 65.26 0.553 / 0.257 / 52.57
DCDP 0.642 / 0.333 / 92.92 0.757 / 0.167 / 42.32 0.719 / 0.282 / 82.23 0.508 / 0.384 / 115.16 0.803 / 0.160 / 44.79
DPS 0.591 / 0.357 / 81.05 0.727 / 0.259 / 75.56 0.647 / 0.285 / 73.33 0.588 / 0.327 / 76.82 0.648 / 0.281 / 74.27 0.693 / 0.284 / 77.61
DAPS-100 0.721 / 0.233 / 69.28 0.716 / 0.194 / 52.89 0.731 / 0.220 / 63.35 0.766 / 0.182 / 54.76 0.685 / 0.239 / 68.11 0.819 / 0.184 / 44.86
DAPS-1K 0.782 / 0.192 / 55.46 0.747 / 0.176 / 50.11 0.786 / 0.179 / 52.73 0.836 / 0.137 / 38.38 0.762 / 0.191 / 57.80 0.839 / 0.163 / 41.09
DAPS++-50 (ours) 0.781 / 0.176 / 45.99 0.812 / 0.141 / 42.14 0.784 / 0.171 / 51.13 0.829 / 0.136 / 37.86 0.745 / 0.194 / 54.86 0.834 / 0.169 / 42.30
@@
Method SR ×4 Inpainting Gaussian Blur Motion Blur Nonlinear Blur HDR
[…]
DDNM 0.487 / 0.458 / 215.27 0.656 / 0.289 / 157.85 0.650 / 0.231 / 90.83
DDRM 0.652 / 0.296 / 111.20 0.602 / 0.336 / 163.67 0.584 / 0.361 / 161.01
DiffPIR 0.443 / 0.370 / 106.75 0.500 / 0.390 / 149.22 0.395 / 0.428 / 154.58 0.152 / 0.683 / 264.79
DPS 0.454 / 0.474 / 208.41 0.623 / 0.343 / 157.49 0.526 / 0.355 / 155.61 0.542 / 0.365 / 161.40 0.520/ 0.365 / 151.53 0.349 / 0.552 / 213.76
DAPS-100 0.610 / 0.309 / 114.38 0.693 / 0.248 / 123.64 0.619 / 0.296 / 117.00 0.710 / 0.209 / 62.03 0.650 / 0.263 / 105.73 0.810 / 0.189 / 43.17
DAPS-1K 0.638 / 0.295 / 109.13 0.715 / 0.229 / 113.96 0.658 / 0.268 / 107.58 0.769 / 0.175 / 47.09 0.720 / 0.212 / 75.94 0.824 / 0.171 / 44.65
DAPS++-50 (ours) 0.653 / 0.283 / 111.72 0.760 / 0.208 / 113.28 0.666 / 0.269 / 117.98 0.754 / 0.185 / 57.40 0.686 / 0.234 / 85.51 0.807 / 0.189 / 46.92
DAPS++-100 (ours) 0.661 / 0.276 / 114.15 0.771 / 0.195 / 105.31 0.663 / 0.273 / 122.88 0.763 / 0.179 / 56.05 0.718 / 0.211 / 75.68 0.807 / 0.191 / 48.49
[…]
@@
Main results. Tab.1 and Tab.2 present the quantitative comparisons against various baselines, with qualitative examples shown in Fig. 3. Across both linear and nonlinear inverse problems, our method achieves competitive—or superior—reconstruction quality with higher SSIM and lower LPIPS and FID, while requiring substantially fewer NFEs than DAPS-1K. Under identical NFE budgets, DAPS++ consistently attains higher fidelity, demonstrating the effectiveness of fully decoupling the prior and likelihood stages and avoiding overfitting in either stage. DAPS++ also shows strong robustness under heavily degraded conditions, such as bicubic $8\times$ super-resolution, where it can still recover meaningful structural information using only the ImageNet pre-trained model.
@@
During the E–M updates, the optimization stages are deliberately non-overlapping. Conventional MAP-style updates in the M-step typically overfit measurement noise unless carefully regularized. In contrast, DAPS++ employs larger step sizes and fewer refinement iterations, leveraging the structure provided by the diffusion-based initialization while avoiding noise amplification. This design yields stable convergence and improved reconstruction quality at lower computational cost.
@@
Small-step ODE solvers introduce notable discretization error when only a limited number of updates are performed. For example, DAPS-1K uses five Euler steps, resulting in an $\mathcal{O}(\Delta t)$ error that produces a weaker initialization compared with Tweedie-based estimation. As a consequence, more MCMC iterations are needed to return to the posterior manifold. DAPS++, by directly predicting the denoised mean in the E-step, avoids this discretization bottleneck and converges more rapidly with significantly fewer total steps. All baseline methods are evaluated on images normalized to $[0,1]$ to ensure consistency and comparability across datasets.
@@
Noise Tolerance. DAPS++ exhibits strong robustness under increasing noise levels. As shown in Tab. 3 and Fig. 4, when $\sigma_{\min}=0.1$ and $\gamma=0.1$, DAPS tends to overfit the noise, producing visible artifacts. In contrast, DAPS++ preserves structural fidelity and perceptual quality even as the noise level increases. As $\gamma$ rises from 0.05 to 0.4, DAPS++ consistently exhibits stronger robustness, especially when $\sigma_{\min}$ is matched to $\gamma$, demonstrating its ability to prevent overfitting. This robustness arises from the EM-style re-noising mechanism and the reduced number of refinement steps, which together prevent overfitting and maintain diffusion-driven variability.
@@ 4.3 Ablation Studies
Efficiency comparison. We assess the efficiency of DAPS++ under varying computational budgets by evaluating configurations with diffusion NFEs ranging from 20 to 200, and comparing them against DAPS variants using 50 to 1K NFEs as well as several representative baselines. All methods are evaluated in terms of single-image sampling time on an NVIDIA A100 (80GB PCIe) GPU and reconstruction quality using the LPIPS metric. As shown in Fig. 5, the trade-off between sampling cost and reconstruction quality consistently favors DAPS++, which achieves higher fidelity at substantially lower computational cost across both linear and nonlinear inverse problems. Even at extremely low computational settings, such as DAPS++–20, the method converges rapidly and produces competitive reconstructions for linear degradation models, suggesting that the fully decoupled formulation provides an effective path toward stable likelihood-driven refinement without relying on heavy diffusion sampling. As computational budgets increase, DAPS++ exhibits progressively improved performance while maintaining a significantly lower cost profile than diffusion-based alternatives that interleave prior and likelihood updates at every timestep. Overall, these results demonstrate that DAPS++ achieves a favorable balance between accuracy and efficiency, making it a scalable and practical solution for real-time or resource-limited inverse imaging applications.
@@
DAPS++-50 (ours) 0.784 / 0.171 0.728 / 0.191 0.674 / 0.229 0.579 / 0.303
Table 3: Quantitative comparison of DAPS-1K and DAPS++-50 under varying noise levels on the FFHQ Gaussian blur task with different $\sigma_{\min}$ is matched to $\gamma$.
Figure 5: Evaluation of sampling time versus reconstruction quality. The x-axis denotes the per-image sampling time on an NVIDIA A100 (80GB PCIe) GPU, and the y-axis shows the LPIPS metric. Results are averaged over 100 FFHQ images for both the Gaussian blur and nonlinear blur tasks.
Hyperparameter choice. Two hyperparameters play a central role in DAPS++: the threshold noise level $\bar{\sigma}$ and the polynomial scheduler exponent. Both parameters directly influence reconstruction quality, particularly when the number of sampling steps is limited and the interaction between diffusion initialization and likelihood refinement becomes more sensitive. The threshold $\bar{\sigma}$ determines when Tweedie-based estimation transitions to a higher-order ODE solver within the diffusion process, allowing efficient denoising at large noise levels and producing a reliable initialization with only a few high-order ODE evaluations. However, even under small noise settings, purely diffusion-generated estimates without data-driven refinement will drift away from the measurement to the next E-step. Therefore, after the higher-order solver step, we still apply an M-step update at the end of the annealing schedule to correct this deviation.
The polynomial scheduler exponent governs the decay rate of the noise levels during annealing and becomes especially important in low-step regimes (e.g., 20–50 steps), as it controls the trade-off between convergence stability and refinement strength. A slower decay stabilizes the M-step updates but introduces more high-noise transitions, which may reduce the effectiveness of likelihood refinement. A faster decay generally accelerates convergence and is often preferred in practice, provided that the M-step remains numerically stable. A comprehensive sensitivity analysis for these hyperparameters, along with practical tuning guidelines, is provided in the Appendix.

### v2

@@ 3.4 Connections to Existing Diffusion-Based Solvers
The two-stage formulation clarifies why DAPS++ achieves its efficiency gains through structural decoupling rather than parameter reduction. Although many diffusion-based inverse problem approaches are presented as posterior samplers, recent studies show that their updates often resemble MAP-style iterative refinements [xu2025rethinking]. The distinction between methods lies in how tightly they couple the diffusion prior with the likelihood at each step, and what computational cost that coupling incurs.
In DPS [chung2022diffusion], the data-consistency term requires differentiating the measurement residual $\|\mathbf{y}-\mathcal{A}(\hat{\mathbf{x}}_{0}(\mathbf{x}_{t}))\|_{2}^{2}$ through Tweedie’s estimator $\hat{\mathbf{x}}_{0}$, propagating gradients back through the score function at every diffusion timestep. This Jacobian-vector product through the score function at each of $T$ steps is the primary computational bottleneck of DPS, and makes it structurally incompatible with higher-order ODE solvers, where additional score evaluations per step would each require a corresponding backward pass. From Eq.˜3, the prior and likelihood gradients are statistically independent, so the DPS update decomposes as
$\mathbf{x}_{t-1}^{\prime}$ $=\underbrace{\mathbf{x}_{t}+\sigma_{t}^{2}\,s_{\theta}(\mathbf{x}_{t},\sigma_{t})+\sigma_{t-1}\mathbf{z}}_{\text{initialization}}$ (10)
$\quad-\underbrace{\eta_{t}\Big[\nabla_{\mathbf{x}_{0}}\,\|\mathbf{y}-\mathcal{A}(\mathbf{x}_{0})\|_{2}^{2}\Big]_{\mathbf{x}_{0}=\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})}}_{\text{refinement}},$
where $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\eta_{t}$ denotes the step size. DPS therefore performs initialization and a single likelihood correction at every timestep, but couples them so that the score network and measurement operator must be evaluated jointly at each step.
DAPS [zhang2025improving] breaks this coupling at the algorithmic level by alternating Tweedie estimation with Langevin dynamics. By applying this directly to $\hat{\mathbf{x}}_{0}$ in image space, the method iteratively updates the clean-image estimate without backpropagating through the score function. However, because DAPS re-invokes the score network at each annealing level to preserve the time-marginal distribution $p(\mathbf{x}_{t}|\mathbf{y})$, the diffusion prior remains an active ingredient throughout sampling. Furthermore, many Langevin steps per annealing level are required to approximately sample from $p(\mathbf{x}_{0}|\mathbf{x}_{t},\mathbf{y})$, which is the necessary condition for theoretically preserving the time-marginal $p(\mathbf{x}_{t}|\mathbf{y})$ at each level. Without sufficiently converged Langevin samples at each stage, this guarantee breaks down.
DAPS++ takes this separation to its logical conclusion: the score network is used only in Stage 1, and Stage 2 operates entirely under the likelihood with no further access to the prior, removing the time-marginal constraint and yielding the approximately 90% reduction in NFEs reported in Sec. 4.
@@ 4 Experiments
Figure 3: (a) Gaussian blur reconstruction on the FFHQ-256 dataset compared with selected baselines. (b) Nonlinear blur reconstruction on ImageNet, comparing DAPS++-50 and DAPS++-100 alongside selected baselines. (c) High dynamic range (HDR) reconstruction on the FFHQ-256 dataset. (d) Motion blur reconstruction on the FFHQ-256 dataset. (e) Box inpainting results on ImageNet, where DAPS++ accurately restores structural details. (f) $8\times$ super-resolution using a pre-trained ImageNet diffusion model, demonstrating strong visual fidelity.
@@ 4.1 Experimental Setup
We evaluate our method using pixel-space pre-trained diffusion models on the FFHQ [chung2022diffusion] and ImageNet [dhariwal2021diffusion] datasets. For the noise schedule, we adopt the POLY(-7) discretization from EDM [karras2022elucidating]. To improve efficiency, $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})$ is generated using Eq.˜7 while $\sigma_{t}>\bar{\sigma}$, and a single RK4 update is used once $\sigma_{t}\leq\bar{\sigma}$. Empirically, we set $\bar{\sigma}=10\times\gamma$ to balance computational cost and reconstruction accuracy. The total number of neural function evaluations (NFEs) remains low because each iteration requires only one network evaluation, further reduced by an annealed step-size schedule.
We use DAPS++-50 (50 annealing steps) for FFHQ and DAPS++-100 (100 steps) for ImageNet. Each DAPS++ iteration performs 4–8 MCMC refinement steps, except for nonlinear deblurring, which requires around 50 refinement steps in the refinement stage. Learning rates are tuned individually for each task. Additional details regarding model configurations, samplers, and hyperparameters appear in the Appendix, along with a discussion of sampling efficiency in Fig.˜5.
@@ 4.2 Results
Method SR $\times$4 Inpainting Gaussian Blur Motion Blur Nonlinear Blur HDR
SSIM / LPIPS / FID SSIM / LPIPS / FID SSIM / LPIPS / FID SSIM / LPIPS / FID SSIM / LPIPS / FID SSIM / LPIPS / FID
DDNM 0.720 / 0.290 / 152.9 0.810 / 0.162 / 94.7 0.804 / 0.216 / 57.8
DDRM 0.820 / 0.191 / 76.1 0.792 / 0.210 / 88.9 0.786 / 0.218 / 89.4
DiffPIR 0.545 / 0.269 / 60.0 0.525 / 0.344 / 63.8 0.514 / 0.293 / 65.3 0.553 / 0.257 / 52.6
DCDP 0.642 / 0.333 / 92.9 0.757 / 0.167 / 42.3 0.719 / 0.282 / 82.2 0.508 / 0.384 / 115.2 0.803 / 0.160 / 44.8
DPS 0.591 / 0.357 / 81.1 0.727 / 0.259 / 75.6 0.647 / 0.285 / 73.3 0.588 / 0.327 / 76.8 0.648 / 0.281 / 74.3 0.693 / 0.284 / 77.6
DAPS-100 0.721 / 0.233 / 69.3 0.716 / 0.194 / 52.9 0.731 / 0.220 / 63.4 0.766 / 0.182 / 54.8 0.685 / 0.239 / 68.1 0.819 / 0.184 / 44.9
DAPS-1K 0.782 / 0.192 / 55.5 0.747 / 0.176 / 50.1 0.786 / 0.179 / 52.7 0.836 / 0.137 / 38.4 0.762 / 0.191 / 57.8 0.839 / 0.163 / 41.1
DAPS++-50 0.781 / 0.176 / 46.0 0.812 / 0.141 / 42.1 0.784 / 0.171 / 51.1 0.829 / 0.136 / 37.9 0.745 / 0.194 / 54.9 0.834 / 0.169 / 42.3
@@
Method SR $\times$4 Inpainting Gaussian Blur Motion Blur Nonlinear Blur HDR
SSIM / LPIPS / FID SSIM / LPIPS / FID SSIM / LPIPS / FID SSIM / LPIPS / FID SSIM / LPIPS / FID SSIM / LPIPS / FID
DDNM 0.487 / 0.458 / 215.3 0.656 / 0.289 / 157.9 0.650 / 0.231 / 90.8
DDRM 0.652 / 0.296 / 111.2 0.602 / 0.336 / 163.7 0.584 / 0.361 / 161.0
DiffPIR 0.443 / 0.370 / 106.8 0.500 / 0.390 / 149.2 0.395 / 0.428 / 154.6 0.152 / 0.683 / 264.8
DPS 0.454 / 0.474 / 208.4 0.623 / 0.343 / 157.5 0.526 / 0.355 / 155.6 0.542 / 0.365 / 161.4 0.520 / 0.365 / 151.5 0.349 / 0.552 / 213.8
DAPS-100 0.610 / 0.309 / 114.4 0.693 / 0.248 / 123.6 0.619 / 0.296 / 117.0 0.710 / 0.209 / 62.0 0.650 / 0.263 / 105.7 0.810 / 0.189 / 43.2
DAPS-1K 0.638 / 0.295 / 109.1 0.715 / 0.229 / 114.0 0.658 / 0.268 / 107.6 0.769 / 0.175 / 47.1 0.720 / 0.212 / 75.9 0.824 / 0.171 / 44.7
DAPS++-50 0.653 / 0.283 / 111.7 0.760 / 0.208 / 113.3 0.666 / 0.269 / 118.0 0.754 / 0.185 / 57.4 0.686 / 0.234 / 85.5 0.807 / 0.189 / 46.9
DAPS++-100 0.661 / 0.276 / 114.2 0.771 / 0.195 / 105.3 0.663 / 0.273 / 122.9 0.763 / 0.179 / 56.1 0.718 / 0.211 / 75.7 0.807 / 0.191 / 48.5
@@
Main results. Tabs.˜1 and 2 present the quantitative comparisons against various baselines, with qualitative examples shown in Fig.˜3. Across both linear and nonlinear inverse problems, our method achieves competitive—or superior—reconstruction quality with higher SSIM and lower LPIPS and FID, while requiring substantially fewer NFEs than DAPS-1K. Under identical NFE budgets, DAPS++ consistently attains higher fidelity, demonstrating the effectiveness of fully decoupling the prior and likelihood stages and avoiding overfitting in either stage. DAPS++ also shows strong robustness under heavily degraded conditions, such as bicubic $8\times$ super-resolution, where it can still recover meaningful structural information using only the ImageNet pre-trained model.
@@
During the alternating initialization-refinement cycles, the optimization stages are deliberately non-overlapping. Conventional MAP-style updates in the refinement stage typically overfit measurement noise unless carefully regularized. In contrast, DAPS++ employs larger step sizes and fewer refinement iterations, leveraging the structure provided by the diffusion-based initialization while avoiding noise amplification. This design yields stable convergence and improved reconstruction quality at lower computational cost.
@@
Small-step ODE solvers introduce notable discretization error when only a limited number of updates are performed. For example, DAPS-1K uses five Euler steps, resulting in an $\mathcal{O}(\Delta t)$ error that produces a weaker initialization compared with Tweedie-based estimation. As a consequence, more MCMC iterations are needed to return to the posterior manifold. DAPS++, by directly predicting the denoised mean in the initialization stage, avoids this discretization bottleneck and converges more rapidly with significantly fewer total steps. All baseline methods are evaluated on images normalized to $[0,1]$ to ensure consistency and comparability across datasets.
@@
Noise Tolerance. DAPS++ exhibits strong robustness under increasing noise levels. As shown in Tabs.˜3 and 4, when $\sigma_{\min}=0.1$ and $\gamma=0.1$, DAPS tends to overfit the noise, producing visible artifacts. In contrast, DAPS++ preserves structural fidelity and perceptual quality even as the noise level increases. As $\gamma$ rises from 0.05 to 0.4, DAPS++ consistently exhibits stronger robustness, especially when $\sigma_{\min}$ is matched to $\gamma$, demonstrating its ability to prevent overfitting. This robustness arises from the structured re-noising mechanism and the reduced number of refinement steps, which together prevent overfitting and maintain diffusion-driven variability.
@@
Figure 5: Sampling time vs. reconstruction quality on an NVIDIA A100 (80GB PCIe) GPU. The y-axis shows LPIPS; results are averaged over 100 FFHQ images for Gaussian blur and nonlinear blur.
@@ 4.3 Ablation Studies
Efficiency comparison. We assess the efficiency of DAPS++ under varying computational budgets by evaluating configurations with diffusion NFEs ranging from 20 to 200, and comparing them against DAPS variants using 50 to 1K NFEs as well as several representative baselines. All methods are evaluated in terms of single-image sampling time on an NVIDIA A100 (80GB PCIe) GPU and reconstruction quality using the LPIPS metric. As shown in Fig.˜5, the trade-off between sampling cost and reconstruction quality consistently favors DAPS++, which achieves higher fidelity at substantially lower computational cost across both linear and nonlinear inverse problems. Even at extremely low computational settings, such as DAPS++–20, the method converges rapidly and produces competitive reconstructions for linear degradation models, suggesting that the fully decoupled formulation provides an effective path toward stable likelihood-driven refinement without relying on heavy diffusion sampling. As computational budgets increase, DAPS++ exhibits progressively improved performance while maintaining a significantly lower cost profile than diffusion-based alternatives that interleave prior and likelihood updates at every timestep. Overall, these results demonstrate that DAPS++ achieves a favorable balance between accuracy and efficiency, making it a scalable and practical solution for real-time or resource-limited inverse imaging applications.
Table 3: Comparison of DAPS-1K and DAPS++-50 under varying noise levels on the FFHQ Gaussian blur task, where $\sigma_{\min}$ is matched to $\gamma$.
@@
DAPS++-50 0.784 / 0.171 0.728 / 0.191 0.674 / 0.229 0.579 / 0.303
Hyperparameter choice. Two hyperparameters play a central role in our method: the threshold noise level $\bar{\sigma}$ and the polynomial scheduler exponent. Both parameters directly influence reconstruction quality, particularly when the number of sampling steps is limited and the interaction between diffusion initialization and likelihood refinement becomes more sensitive. The threshold $\bar{\sigma}$ determines when Tweedie-based estimation transitions to a higher-order ODE solver within the diffusion process, allowing efficient denoising at large noise levels while producing a reliable initialization with only a few high-order ODE evaluations. However, even under small noise settings, purely diffusion-generated estimates without data-driven refinement will drift away from the measurement; therefore, after the higher-order solver step, we still apply a refinement update at the end of the annealing schedule to correct this deviation.
The polynomial scheduler exponent governs the decay rate of the noise levels during annealing and becomes especially important in low-step regimes (e.g., 20–50 steps), as it controls the trade-off between convergence stability and refinement strength. A slower decay stabilizes the refinement updates but introduces more high-noise transitions, which may reduce the effectiveness of likelihood refinement. A faster decay generally accelerates convergence and is often preferred in practice, provided that the refinement stage remains numerically stable. A comprehensive sensitivity analysis for these hyperparameters, along with practical tuning guidelines, is provided in the Appendix.